OpenStudy (anonymous):

A philatelic collection is comprised of stamps from Germany, US, and India. . If the ratio of american stamps to Indian stamps is 5 to 2. The ratio of German stamps to Indian stamps is 5 to 1. What is the ratio american to German stamps?

OpenStudy (anonymous):

please show the work

OpenStudy (anonymous):

US : Germany : India = 5 : 10 : 2 US : Germany = 5 : 10 = 1 : 2

OpenStudy (moongazer):

i also have a problem that is similar with this i also want to learn how this works :)

OpenStudy (anonymous):

call america A, Germany G, india I

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

A:I = 5:2 G:I = 5:1

OpenStudy (anonymous):

where : means "to" , as in, ratio, like 1/2 is 1:2 ,

OpenStudy (anonymous):

ok got you

OpenStudy (anonymous):

we want to find A:G or A/ G

OpenStudy (anonymous):

A:G = something : something

OpenStudy (anonymous):

we know that A/I = 5/2, and G/I = 5/1, so lets solve for A and G in terms of I , A = 5/2* I , and G = 5/1 * I

OpenStudy (anonymous):

so A/G = ( 5/2) / ( 5/1) = (5/2) / ( 5) = (5/2) * 1/5 = 5/10 = 1/2

OpenStudy (anonymous):

oh yeah, i forgot to mention, the I's cancel when you do A/G

OpenStudy (anonymous):

leaveing you with just numbers

OpenStudy (anonymous):

A/G = ( 5/2*I ) / ( 5/1* I ) I means the letter I

OpenStudy (anonymous):

yeah sure. Awesome man, I finally got it :)

OpenStudy (anonymous):

oh that looks confusing sorry

OpenStudy (anonymous):

ok lets use lower case i , so

OpenStudy (anonymous):

no its ok. I got it

OpenStudy (anonymous):

A/G = ( 5/2*i ) / ( 5/1* i ) , and the i's cancel

OpenStudy (moongazer):

could it be 5/2=5/1? so it will be 5/10 or 1:2?

OpenStudy (anonymous):

right, 1:2

OpenStudy (anonymous):

you have a typo, should be 5/2 : 5/1 , not

OpenStudy (moongazer):

oh yes

OpenStudy (moongazer):

so it is correct?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

cantor - thats a good explanation of the process. Its peculiar - though i consider myself to be good at math, i've always had a mental block when it comes to ratios - thank you.

OpenStudy (anonymous):

no problem , :)

OpenStudy (anonymous):

heres a shorter proof

OpenStudy (anonymous):

A/ i = 5/2, G/i = 5/1, and notice that if i do ( A/i) / ( G/i ) , the i's will cancel, so (A/i) / ( G/i) =( 5/2) / ( 5/1) , but the left side cancels to A/G, and the right side is 1:2 after simplification

OpenStudy (anonymous):

That is very clear, Cantor! Thanks for that!